OpenStudy (anonymous):

Simplify (5/4x^2y)-(y/14xz)

OpenStudy (anonymous):

\[\frac {5}{4x^{2}y}-\frac {y}{14xz}\] Take LCM \[\frac {5*(14)xz-4*4x^{2y}}{4x^{2}y*14xz}\] \[\frac {60xz-16x^{2}y}{4x^{2}y*14xz}\] Taking commons, \[\frac {4x(15z-4xy)}{2x(2xy*7z)}\] \[\frac {2(15z-4xy)}{(2xy*7z)}\]

OpenStudy (anonymous):

Thank you once again :)

OpenStudy (anonymous):

welcome